OpenStudy (anonymous):

a lab technician mixed a 510 ml solution of water and alcohol. if 3% of the solution is alcohol, how many milliliter of a alcohol were used, and how many milliliters of water were used?

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (xapproachesinfinity):

let x and y be the amount of alcohol and water in the solution respectively then we have x+y=510 ml if 3% of the solution is alcohol then 97% of the solution is water so 0.03x+0.97y=510 we solve the system \(\large \rm \begin{cases} x+y=510 \\ 0.03x+0.97y=510 \\ \end{cases}\)